OpenStudy (gabbyalicorn):

Forty-two is what percent of 70? A. 21% B. 60% C. 74% D. 167% 21?

OpenStudy (anonymous):

42/70 = 6/10 = 60/100 = 60%

OpenStudy (anonymous):

aka, B

OpenStudy (gabbyalicorn):

ty

OpenStudy (anonymous):

your welcome :)